OpenStudy (anonymous):

Find the exact value of y 1.) y=tan^-1 (tan 3pi/4) 2.) y=cos^-1 (cos7pi/6) 3.) y=cos(sin^-1 1/2)

OpenStudy (amistre64):

a simple solution is to draw out the triangles

OpenStudy (amistre64):

the inverses cancel out each other when they are the same base function; but the angle changes to fit the range of the inverse

OpenStudy (amistre64):

if i recall correctly tan and cos inverses are on top of the circle (0 to pi); and sin is on the side (-pi/2 to pi/2)

OpenStudy (callisto):

Actually, they are all special angles... 1.) y=tan^-1 (tan 3pi/4) = 1=tan^-1 (-tan (pi/4)) = tan^-1 (-1) = 3pi/4 or7pi/4 2.) y=cos^-1 (cos7pi/6) = cos^-1 (-cos pi/6) = 5pi/6 or 7pi/6 3.) y=cos(sin^-1 1/2) = cos (pi/6) or cos (5pi/6) = sqrt(3) / 2 or - sqrt(3) / 2

OpenStudy (amistre64):

the inverses are one to one, so they dont give 2 different angles

OpenStudy (callisto):

@amistre64 I didn't know that... Sorry!!! and thanks too !

OpenStudy (anonymous):

1) tan^-1(tan(3pi/4) tan^-1(-1) = -pi/4

OpenStudy (anonymous):

2) cos^-1 ( cos (7pi /6)) = cos^-1 ( cos (pi + pi/6)) = cos^-1 ( - cos (pi/6)) = cos^-1 ( - sqrt(3)/2) = 5pi/6 Now try rest similarly

OpenStudy (anonymous):

Just remember See the attachement

OpenStudy (amistre64):